OpenStudy (anonymous):

How many solutions are there for the system shown below? x^2+y^2=25 x-y^2=-5 A. 1 B. 2 C. 4 D. 3

OpenStudy (shamil98):

y^2 = x + 5 x^2 + x + 5 = 25 x^2 + x - 20 = 0 Can you do it from here?

OpenStudy (anonymous):

would that be 2 solutions because it would be (x+10)(x-10)?

OpenStudy (anonymous):

@shamil98